OpenStudy (anonymous):

sole:logx-log4=log12

OpenStudy (anonymous):

logx-log4 = log12 log(x/4)=log12 x/4 = 12 x = 12·4 = 48

OpenStudy (anonymous):

what??? it's that easy???

OpenStudy (anonymous):

Yeah, you have to know logarithm properties.

OpenStudy (anonymous):

soo..it's just 48? not logx48?

OpenStudy (anonymous):

loga + logb = log ab loga - logb = log a/b

OpenStudy (anonymous):

yeah, x = 48

OpenStudy (anonymous):

may i ask one more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

solve: logx64=3 ....the x is like below the log

OpenStudy (anonymous):

172?

OpenStudy (anonymous):

Convert it into exponent form. Do you know how to?

OpenStudy (anonymous):

x^3=64?

OpenStudy (anonymous):

Yes. x³ = 64 x = \(\sqrt[3]{64}\) = 8

OpenStudy (anonymous):

thank you so much .....uhhhhmmm...is it rainbow dash?

OpenStudy (anonymous):

My bad, it's 4, not 8. Yes, it is rainbow dash lol.

OpenStudy (anonymous):

RAINBOW DASH!!! you got it wrong >:3

OpenStudy (anonymous):

lol thank you so much :)

OpenStudy (anonymous):

lol glad I helped.